# ContrastAlign: Toward Robust BEV Feature Alignment via Contrastive Learning for Multi-Modal 3D Object Detection

### v2

@@ Abstract
In the field of 3D object detection tasks, fusing heterogeneous features from LiDAR and camera sensors into a unified Bird’s Eye View (BEV) representation is a widely adopted paradigm. However, existing methods are often compromised by imprecise sensor calibration, resulting in feature misalignment in LiDAR-camera BEV fusion. Moreover, such inaccuracies result in errors in depth estimation for the camera branch, ultimately causing misalignment between LiDAR and camera BEV features. In this work, we propose a novel ContrastAlign approach that utilizes contrastive learning to enhance the alignment of heterogeneous modalities, thereby improving the robustness of the fusion process. Specifically, our approach includes the L-Instance module, which directly outputs LiDAR instance features within LiDAR BEV features. Then, we introduce the C-Instance module, which predicts camera instance features through RoI (Region of Interest) pooling on the camera BEV features. We propose the InstanceFusion module, which utilizes contrastive learning to generate similar instance features across heterogeneous modalities. We then use graph matching to calculate the similarity between the neighboring camera instance features and the similarity instance features to complete the alignment of instance features. Our method achieves state-of-the-art performance, with an mAP of 70.3%, surpassing BEVFusion by 1.8% on the nuScenes validation set. Importantly, our method outperforms BEVFusion by 7.3% under conditions with misalignment noise.
1 Introduction
3D object detection is one of the fundamental tasks for achieving reliable environment perception in autonomous driving tasks [44, 70]. The task aims to accurately identify and locate obstacles such as cars and pedestrians, to provide precise and real-time data that informs the autonomous driving system and reach correct driving decisions. The current standard for achieving safe, robust, and high-precision detection may require the fusion of heterogeneous modalities: Due to the inherent sparsity of point clouds, LiDAR-only methods [23, 62, 71, 12, 41, 35, 40, 55, 46] struggle to detect small or distant objects, making them insufficient for robust 3D detection. In contrast, these objects remain clear and distinguishable in high-resolution images, containing rich semantic representations [56]. The complementary roles of point clouds and images have prompted researchers to leverage the strengths of heterogeneous modalities to design detectors.
Based on different fusion strategies, heterogeneous 3D object detection can be mainly categorized into point-level [42, 52, 18, 33, 53], feature-level [1, 27, 64, 10, 45, 43, 49, 48] methods and the currently dominant BEV-based methods [34, 29]. BEV-based methods integrate LiDAR and camera modalities into a shared BEV representation space. Although BEV-based methods [34, 29] have demonstrated promising performance, they still suffer BEV feature misalignment. There are two main causes of this problem, including sensor misalignment (see Figure 1 (a)) and depth inaccuracy (see Figure 1 (b)). Firstly, for sensor misalignment, as pointed out in BEVDepth [28], GraphBEV [47], ObjectFusion [4], calibration matrix errors between LiDAR and camera sensors can result in feature misalignment. Second, their fusion process relies heavily on the accuracy of depth estimation from camera to BEV (e.g., LSS[38]), with depth inaccuracy leading to feature misalignment further.
@@ 1 Introduction
Figure 1: Motivation of ContrastAlign. (a) Misalignment persists in real-world scenarios, particularly during vibrations while driving on bumpy roads. Since such noise occurs frequently, it cannot be easily mitigated through online calibration. (b) During the fusion of LiDAR BEV and camera BEV in BEVFusion [34, 29], the errors in depth within Camera-to-BEV modules (e.g., LSS [38]) lead to misalignment of BEV features. (c) We propose a novel ContrastAlign framework, which achieves BEV feature alignment by leveraging contrastive learning to calculate the similarity between heterogeneous instance features of LiDAR and camera. Empirical results reveal that our ContrastAlign surpasses the BEVFusion [34] by a margin of 1.8% mAP on the nuScenes validation dataset [3] and by over 7.3% in the setting with noisy misalignment [13].
@@
The issue of feature alignment exists not only in multi-modal 3D object detection [52, 53, 34, 29, 48] but also in multi-modal tasks involving text and images [68, 60, 56, 44]. With the development of multi-modal foundational models, more researchers [39, 57, 20, 21, 5] are paying attention to heterogeneous modality alignment for the purpose of modality consistency. Contrastive alignment [20, 68, 60] can be seen as a mutually refining process where each modality contributes to and benefits from the exchange of knowledge. Through contrastive learning, the iterative interaction between heterogeneous modalities continuously evolves, enhancing their ability for general understanding within a single model and addressing the alignment issues of heterogeneous modalities [20]. Therefore, inspired by the above idea, we apply its concept to the task of 3D object detection to solve the feature misalignment between LiDAR and camera BEV features [34, 29] .
In this work, we propose ContrastAlign, which leverages contrastive learning to enhance the alignment of heterogeneous modalities, thereby improving the robustness of LiDAR-camera BEV feature fusion as shown in Figure 1 (c). Specifically, we propose the L-Instance module, which directly outputs LiDAR instance features within LiDAR BEV features. Then, we introduce the C-Instance module, which predicts camera instance features through RoI (Region of Interest) Pooling on camera BEV features. The LiDAR instance features are then projected onto image instance features, and contrastive learning is employed to generate similar instance features between LiDAR and camera. Subsequently, through graph matching, neighboring camera instance features are matched to calculate similarity and construct positive and negative samples. During inference, the aligned features with high similarity in neighboring camera instance features are selected as alignment features to achieve BEV feature alignment. Extensive experiments have demonstrated the effectiveness of our ContrastAlign, with significant performance improvement on the nuScenes [3] dataset, especially at the misaligned noisy setting [13].
2 Related Work
2.1 Multi-modal 3D Object Detection
@@ 2.2 Contrastive Learning
3 Method
To address the feature misalignment issue, we propose a robust fusion framework named ContrastAlign. The overview of our framework is provided in Figure 2. In the following sections, we first introduce the general overview of the proposed ContrastAlign in Section 3.1. Subsequently, we delve into the details of the C-Instance and L-Instance modules in Section 3.2. After that, in Section 3.3, we elaborate on the crucial design steps of the InstanceFusion module.
3.1 Overall Framework of ConstrastAlign
@@
C-Instance and L-Instance Modules. Previous methods like BEVFusion [34, 29] directly concatenate $F^{C}_{B}$ and $F^{L}_{B}$. The operation can not distinct instance features and background features in the BEV space. Compared to background features, instance features such as cars and pedestrians are more important for 3D detection. To obtain LiDAR and camera instance features, we introduce C-Instance and L-Instance modules, which predict instance features through score filtering and ROI Pooling, as detailed in Section 3.2.
InstanceFusion. We introduce the InstanceFusion module to achieve cross-modal alignment during the BEV fusion of LiDAR and camera. This module, detailed in Section 3.3, represents our core innovation. It leverages the C-Instance and L-Instance modules to provide LiDAR and camera instance features, which are then aligned into the BEV space.
Detection Head. We follow TransFusion [1] to generate the final 3D detection results. During the training process, we incorporate the InfoNCE loss as described in Section 3.3. Meanwhile, we utilize Focal loss [30] and L1 loss for classification and 3D bounding box regression, respectively.
3.2 C-Instance and L-Instance Modules
@@
As shown in Figure 3, C-Instance and L-Instance modules aim to extract instance features from the BEV features of LiDAR and camera. For the camera BEV features $F^{C}_{B}$, we follow BEVDet [17] employ CenterPoint Head [66] to generate proposals $P_{C}$. For the LiDAR BEV features $F^{L}_{B}$, we employ VoxelNeXt Head [9] to generate proposals $P_{L}$. During inference processing, sparse max pooling [8] is employed to score the selected proposals. LiDAR and camera proposals, $P_{C}$ and $P_{L}$, consist of features including positions $(\delta_{x},\delta_{y})\in\mathbb{R}^{2}$, heights $z_{c}\in\mathbb{R}$, dimensions $(w,h,l)\in\mathbb{R}^{3}$, orientation angles $\alpha\in\mathbb{R}$, and proposal scores and labels.
@@
Subsequently, we employ feature sampling [66] to generate RoI features $F_{C}^{RoI}$ and $F_{L}^{RoI}$ from $F^{C}_{B}$ and $F^{L}_{B}$. Specifically, for any proposal $p_{c}\in P_{C}$ or $p_{l}\in P_{L}$, we select its positions $(\delta_{x},\delta_{y})$ and sample points at the center of each boundary line of the bounding box. Each proposal generates a set of sampling points $[c,c_{\uparrow},c_{\downarrow},c_{\leftarrow},c_{\rightarrow}]$, where $c$ denotes its positions $(\delta_{x},\delta_{y})$. Apart from $c$, the other sampling points yield sampling results on the LiDAR and camera BEV features using interpolation algorithms [66]. Finally, the LiDAR and camera RoI features, $F_{C}^{RoI}\in\mathbb{R}^{B_{S}\times N\times(5\times C_{RoI})}$ and $F_{L}^{RoI}\in\mathbb{R}^{B_{S}\times M\times(5\times C_{RoI})}$, are formed by concatenating all sampling features, where $B_{S}$ denotes batch size, $N$ and $M$ denote the numbers of LiDAR and camera RoI features, 5 denotes the number of $[c,c_{\uparrow},c_{\downarrow},c_{\leftarrow},c_{\rightarrow}]$, and $C_{RoI}$ denotes the channel of $F_{C}^{RoI}$ and $F_{L}^{RoI}$.
3.3 InstanceFusion
@@
We propose the InstanceFusion module to achieve cross-modal alignment of instance features. For positive samples, we project LiDAR instance features onto the camera BEV space and consider the matched camera instance features as positive samples. For negative samples, we select the $K$ nearest neighboring instance features around the camera positive samples as negative samples. During inference, LiDAR instance features calculate similarity scores with $K$ neighboring camera instances and select the instance feature with the highest score as the matching aligned instance feature. The whole process of InstanceFusion module are as follows.
@@
First, due to the absence of depth errors in the LiDAR BEV feature implementation, we utilize the generated LiDAR RoI features to retrieve the corresponding camera RoI features. For any pair of bounding box predictions from $P_{C}$ with $N$ samples and $P_{L}$ with $M$ samples, we calculate the Intersection over Union (IoU) by computing the intersection and union areas of their 2D bounding boxes. We define the bounding box sets $\text{B}_{C}^{RoI}=\{(x_{i},y_{i},w_{i},h_{i})|i\in[1,2...N]\}$ and $\text{B}_{L}^{RoI}=\{(x_{i},y_{i},w_{i},h_{i})|i\in[1,2...M]\}$ based on the positions $(\delta_{x},\delta_{y})$ and dimensions $(w,h)$ of proposals in Section 3.2. To obtain the groups of positive samples for $F^{pos}$, we execute that
@@
After obtaining the positive samples, we proceed to construct negative samples. Due to the impact of feature misalignment, offsets occur in a region. We utilize the KNN algorithm like KD-Tree to construct the neighbors’ relations of positive samples and consider each positive sample’s neighbor features as negative samples. Specifically, the construction of negative samples is as follows.
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics
We evaluate the 3D object detection performance of the proposed ContrastAlign by comparing it with other state-of-the-art approaches on the nuScenes benchmark [3], which is collected with a 32-beam LiDAR and 6 cameras. The dataset is generally split into 700/150/150 scenes for training/validation/testing. The six images cover a 360-degree surrounding, and the dataset provides calibration matrices that enable precise projection from 3D points to 2D pixels. It requires detecting 10 object categories that are commonly observed in driving scenarios. We use mAP and NDS across all categories as the primary metrics for evaluation following [1, 34, 29]. Note that the NDS metric is a weighted average of mAP and other breakdown metrics (e.g., translation, scale, orientation, velocity, and attributes errors).
4.1.2 Implementation Details
We implement ContrastAlign within PyTorch [37], built upon the open-source BEVFusion [34] and OpenPCDet [50]. In the LiDAR branch, feature encoding is performed by using TransFusion-L [1] to obtain LiDAR BEV features, with voxel dimensions set to [0.075m, 0.075m, 0.2m] and point cloud ranges specified as [-54m, -54m, -5m, 54m, 54m, 3m] along the X, Y, and Z axes, respectively. The camera branch employs a Swin Transformer [32] as the camera backbone, integrating heads of numbers 3, 6, 12, and 24, and utilizes FPN [16] for fusing multi-scale feature maps. The resolution of input images is adjusted and cropped to 256 $\times$ 704. In LSS [38] configuration, frustum ranges are set with X coordinates [-54m, 54m, 0.3m], Y coordinates [-54m, 54m, 0.3m], Z coordinates [-10m, 10m, 20m], and depth ranges are set to [1m, 60m, 0.5m]. During training, we apply data augmentation for 10 epochs, which includes random flips, rotations (within the range [$-\frac{\pi}{4}$, $\frac{\pi}{4}$]), translations (with std=0.5), and scaling in the range of 0.9 to 1.1 for LiDAR data enhancement. We utilize CBGS [72] to resample the training data. Additionally, we use random rotation in [$-5.4^{\circ}$, $5.4^{\circ}$] and random resizing in [0.38, 0.55] to augment the images. The Adam optimizer [22] is used with a one-cycle learning rate policy, setting the maximum learning rate to 0.001 and weight decay to 0.01. The batch size is 24, and training is conducted on 8 NVIDIA GeForce RTX 3090 24G GPUs. During inference, we remove Test Time Augmentation (TTA) data augmentation, and the batch size is set to 1 on an A100 GPU. All latency measurements are taken on the same workstation with an A100 GPU.
Table 1: Comparison with the SOTA methods on the nuScenes validation set under clean setting and noisy misalignment setting. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘ T.C.’ are short for construction vehicles, motorcycles, pedestrians, and traffic cones.
Setting Method Modality mAP NDS Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
Clean TransFusion-L [1] L 65.1 70.1 86.5 59.6 25.4 74.4 42.2 74.1 72.1 56.0 86.6 74.1

### v3

@@ Abstract
In the field of 3D object detection tasks, fusing heterogeneous features from LiDAR and camera sensors into a unified Bird’s Eye View (BEV) representation is a widely adopted paradigm. However, existing methods often suffer from imprecise sensor calibration, leading to feature misalignment in LiDAR-camera BEV fusion. Moreover, such inaccuracies cause errors in depth estimation for the camera branch, aggravating misalignment between LiDAR and camera BEV features. In this work, we propose a novel ContrastAlign approach that utilizes contrastive learning to enhance the alignment of heterogeneous modalities, thereby improving the robustness of the fusion process. Specifically, our approach comprises three key components: (1) the L-Instance module, which extracts LiDAR instance features within the LiDAR BEV features; (2) the C-Instance module, which predicts camera instance features through Region of Interest (RoI) pooling on the camera BEV features; (3) the InstanceFusion module, which employs contrastive learning to generate consistent instance features across heterogeneous modalities. Subsequently, we use graph matching to calculate the similarity between the neighboring camera instance features and the similarity instance features to complete the alignment of instance features. Our method achieves SOTA performance, with an mAP of 71.5%, surpassing GraphBEV by 1.4% on the nuScenes val set. Importantly, our method excels BEVFusion under conditions with spatial & temporal misalignment noise, improving mAP by 1.4% and 11.1% on nuScenes dataset. Notably, on the Argoverse2 dataset, ContrastAlign outperforms GraphBEV by 1.0% in mAP, indicating that the farther the distance, the more severe the feature misalignment and the more effective.
Index Terms:
3D Object Detection, BEV perception, Autonomous Driving, Multi-modal fusion, Feature Misalignment
I Introduction
@@
Figure 1: Motivation of ContrastAlign. (a) Misalignment persists in real-world scenarios, particularly during vibrations while driving on bumpy roads. Since such noise occurs frequently, it cannot be easily mitigated through online calibration. (b) During the fusion of LiDAR BEV and camera BEV in BEVFusion  [1, 2], the errors in depth within Camera-to-BEV modules (e.g., LSS  [3]) lead to misalignment of BEV features. (c) Our contrastAlign framework achieves BEV feature alignment by leveraging contrastive learning to calculate the similarity between heterogeneous instance features of LiDAR and camera. In the BEV fusion stage, the direct concatenation method of BEVFusion may have misalignment issues (gray solid line), but our ContrastAlign corrects this problem through instance alignment.
3D object detection is one of the fundamental tasks for achieving reliable environment perception in autonomous driving tasks [4, 5, 6, 7]. The task aims to accurately identify and locate obstacles such as cars and pedestrians, to provide precise and real-time data that informs the autonomous driving system to reach correct driving decisions [8]. The current standard for achieving safe, robust, and high-precision detection may require the fusion of heterogeneous modalities. Due to the inherent sparsity of point clouds, LiDAR-only methods [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] struggle to detect small or distant objects, making them insufficient for robust 3D detection. In contrast, these objects remain clear and distinguishable in high-resolution images, containing rich semantic representations [22]. The complementary roles of point clouds and images have prompted researchers to leverage the strengths of heterogeneous modalities to design detectors.
Based on different fusion strategies, heterogeneous 3D object detection can be mainly categorized into point-level [23, 24, 25, 26, 27], feature-level [28, 29, 30, 31, 32, 33, 34, 35, 36] methods and the currently dominant BEV-based methods [1, 2]. BEV-based methods integrate LiDAR and camera modalities into a shared BEV representation space. Although BEV-based methods [1, 2] have demonstrated promising performance, they still suffer BEV feature misalignment. There are two main causes of this problem, including sensor misalignment (see Figure 1 (a)) and depth inaccuracy (see Figure 1 (b)). Firstly, for sensor misalignment, as pointed out in BEVDepth [37], GraphBEV [38], ObjectFusion [39], calibration matrix errors between LiDAR and camera sensors can result in feature misalignment. Second, their fusion process relies heavily on the accuracy of depth estimation from camera to BEV (e.g., LSS[3]), with depth inaccuracy leading to feature misalignment further.
@@
The issue of feature alignment exists not only in multi-modal 3D object detection  [24, 27, 1, 2] but also in multi-modal tasks involving text and images  [44, 45]. With the development of multi-modal foundational models, more researchers  [46, 47, 48, 49, 50] are paying attention to heterogeneous modality alignment for the purpose of modality consistency. Contrastive alignment  [48, 44, 45] can be seen as a mutually refining process where each modality contributes to and benefits from the exchange of knowledge. Through contrastive learning, the iterative interaction between heterogeneous modalities continuously evolves, enhancing their ability for general understanding within a single model and addressing the alignment issues of heterogeneous modalities  [48]. Therefore, inspired by the above idea, we apply its concept to the task of 3D object detection to solve the feature misalignment between LiDAR and camera BEV features.
In this work, we propose ContrastAlign, which leverages contrastive learning to enhance the alignment of heterogeneous modalities, thereby improving the robustness of LiDAR-camera BEV feature fusion as shown in Figure 1 (c). Specifically, we propose the L-Instance module, which directly outputs LiDAR instance features within LiDAR BEV features. Then, we introduce the C-Instance module, which predicts camera instance features through RoI (Region of Interest) Pooling on camera BEV features. The LiDAR instance features are then projected onto image instance features, and contrastive learning is employed to generate similar instance features between LiDAR and camera. Subsequently, through graph matching, neighboring camera instance features are matched to calculate similarity and construct positive and negative samples. During inference, the aligned features with high similarity in neighboring camera instance features are selected as alignment features to achieve BEV feature alignment. Extensive experiments have demonstrated the effectiveness of our ContrastAlign, with significant performance improvement on the nuScenes  [51] dataset, especially at the misaligned noisy setting  [52]. This is noteworthy, as we focus on object-level alignment and effectively address temporal misalignment using only single-frame multi-modal data.
II Related Work
II-A Multi-modal 3D Object Detection
@@ II-B Contrastive Learning
III Is Contrastive Learning All You Need for Feature Alignment?
In this section, we conduct a thorough analysis of how contrastive learning addresses the alignment issues between multi-modal inputs (LiDAR and Camera) to substantiate the validity of our approach. The following sections detail its mathematical principles and rationale for feature alignment.
Let $X=\{x_{1},x_{2},\ldots,x_{n}\}$ represent a set of samples, where each sample $x_{i}$ generates a feature representation $f(x_{i})$. A positive pair $(x_{i},x_{j})$ consists of different views of the same object, while a negative pair $(x_{i},x_{k})$ comprises different objects.
The objective of contrastive learning is to minimize the following contrastive loss function:
$L(i,j)=-\log\frac{\exp(\text{sim}(f(x_{i}),f(x_{j}))/\tau)}{\sum_{k=1}^{n}\exp(\text{sim}(f(x_{i}),f(x_{k}))/\tau)}$ (1)
Here, $\text{sim}(a,b)$ denotes the similarity metric, typically using cosine similarity:
$\text{sim}(a,b)=\frac{a\cdot b}{\|a\|\|b\|}$ (2)
The temperature parameter $\tau$ controls the smoothness of the distribution, and $\mathbb{1}_{[k\neq i]}$ ensures negative pairs exclude the current sample.
In the feature alignment process, let $F_{L}$ and $F_{C}$ represent the features extracted from LiDAR and camera modalities, respectively. The goal is to utilize contrastive learning to bring these features closer in a shared representation space. Specifically, for positive pairs $(F_{L}^{i},F_{C}^{j})$, we aim to minimize their distance $d(F_{L}^{i},F_{C}^{j})$, while maximizing the distance for negative pairs $(F_{L}^{i},F_{C}^{k})$:
$d(F_{L}^{i},F_{C}^{j})=\|F_{L}^{i}-F_{C}^{j}\|^{2}$ (3)
This leads to the following objective function:
$\min_{F_{L},F_{C}}\sum_{(i,j)\in P}d(F_{L}^{i},F_{C}^{j})-\sum_{(i,k)\in N}d(F_{L}^{i},F_{C}^{k})$ (4)
where $P$ denotes the set of positive pairs and $N$ the set of negative pairs.
IV Methodology
To address the feature misalignment issue, we propose a robust fusion framework named ContrastAlign which includes our newly designed C-Instance and L-Instance modules, and a InstanceFusion module. The overview of our framework is provided in Figure 2. In the following sub-sections, we first introduce the general overview of the proposed ContrastAlign. Subsequently, we delve into the details of the C-Instance and L-Instance modules. After that, we elaborate on the crucial design steps of the InstanceFusion module.
IV-A Overview of ConstrastAlign
@@
C-Instance and L-Instance Modules. Previous methods like BEVFusion [1, 2] directly concatenate $F^{C}_{B}$ and $F^{L}_{B}$. The operation can not distinct instance features and background features in the BEV space. Compared to background features, instance features such as cars and pedestrians are more important for 3D detection. To obtain LiDAR and camera instance features, we introduce C-Instance and L-Instance modules, which predict instance features through score filtering.
InstanceFusion. We introduce the InstanceFusion module to achieve cross-modal alignment during the BEV fusion of LiDAR and camera. This module, detailed in InstanceFusion Section, represents our core innovation. It leverages the C-Instance and L-Instance modules to provide LiDAR and camera instance features, which are then aligned into the BEV space.
Detection Head. We follow TransFusion  [28] to generate the final 3D detection results. During the training process, we incorporate the InfoNCE loss as described in Eq. (6). Meanwhile, we utilize Focal loss  [66] and L1 loss for classification and 3D bounding box regression, respectively.
IV-B C-Instance and L-Instance Modules
@@
As shown in Figure 3, C-Instance and L-Instance modules aim to extract instance features from the BEV features of LiDAR and camera. Generally, extracting instance features requires point-based operations [69], but these methods are inefficient. We attempt to avoid time-consuming operations like FPS [13], and therefore, choose to use an additional detection head to perform this step on the BEV feature map. For the camera BEV features $F^{C}_{B}$, we follow BEVDet [70], employ CenterPoint Head [67] to generate proposals $P_{C}$. For the LiDAR BEV features $F^{L}_{B}$, we employ VoxelNeXt Head [68] to generate proposals $P_{L}$. It is worth noting that, influenced by submanifold sparse convolution [10, 55], we adopt multiple stacked max pool layers for sparse max pooling during the inference of $P_{C}$, aiming to avoid using NMS. During inference processing, sparse max pooling [55] is employed to score the selected proposals. LiDAR and camera proposals, $P_{C}$ and $P_{L}$, consist of features including positions $(\delta_{x},\delta_{y})\in\mathbb{R}^{2}$, heights $z_{c}\in\mathbb{R}$, dimensions $(w,h,l)\in\mathbb{R}^{3}$, orientation angles $\alpha\in\mathbb{R}$, and proposal scores and labels.
@@
Subsequently, we employ feature sampling [67] to generate RoI features $F_{C}^{RoI}$ and $F_{L}^{RoI}$ from $F^{C}_{B}$ and $F^{L}_{B}$. Specifically, for any proposal $p_{c}\in P_{C}$ or $p_{l}\in P_{L}$, we select its positions $(\delta_{x},\delta_{y})$ and sample points at the center of each boundary line of the bounding box. Each proposal generates a set of sampling points $c,c_{\uparrow},c_{\downarrow},c_{\leftarrow},c_{\rightarrow}]$, where $c$ denotes its positions $(\delta_{x},\delta_{y})$. Apart from $c$, the other sampling points yield sampling results on the LiDAR and camera BEV features using interpolation algorithms [67]. Finally, the LiDAR and camera RoI features, $F_{C}^{RoI}\in\mathbb{R}^{B_{S}\times N\times(5\times C_{RoI})}$ and $F_{L}^{RoI}\in\mathbb{R}^{B_{S}\times M\times(5\times C_{RoI})}$, are formed by concatenating all sampling features, where $B_{S}$ denotes batch size, $N$ and $M$ denote the numbers of LiDAR and camera RoI features, 5 denotes the number of $c,c_{\uparrow},c_{\downarrow},c_{\leftarrow},c_{\rightarrow}]$, and $C_{RoI}$ denotes the channel of $F_{C}^{RoI}$ and $F_{L}^{RoI}$.
IV-C InstanceFusion
Input:
ROI features: $F_{C}^{RoI}$ and $F_{L}^{RoI}$
Position Index: $B_{C}^{RoI}$ and $B_{L}^{RoI}$
IoU threshold: $\tau=0.2$
Number of positive samples’ neighbors: $K=8$
Output:
Fusion features $F^{Fuse}$
1
2while InstanceFusion do
3   Function IoU ($B_{C}^{RoI}$, $B_{L}^{RoI}$)
4     for $i=1$ to NN, $j=1$ to $M$ do
5       $IoU_{i,j}=\frac{A_{i}}{A_{u}}$
6       end for
7
8
9      $I^{pos}=f^{pos}(IoU(B_{C}^{RoI},B_{L}^{RoI}),\tau)$
10       $F^{pos}=f^{F}_{pos}(F_{C}^{RoI},F_{L}^{RoI},I^{pos})$
11       for positions $q$ of $B_{L}^{RoI}$ do
12         Function Shoot ($q$, $K$)
13           $NBR=\text{KD-Tree}(q,K)$ $F^{neg}=f^{F}_{neg}(F_{C}^{RoI},NBR)$
14
15
16           end for
17          Function InfoNCELoss ($F^{pos}$, $F^{neg}$)
18             for each $F^{neg}_{i}$ in $F^{neg}$ do
19               $\mathcal{L}_{align}=-\log\frac{\exp(F^{pos}_{L}\cdot F^{pos}_{C})}{\sum_{k=1}^{K}\exp(F^{pos}_{L}\cdot F^{neg}_{i})}$
20               end for
21
22
23
24               end while
Algorithm 1 InstanceFusion Workflow
@@ 17          Function InfoNCELoss ($F^{pos}$, $F^{neg}$)
We propose the InstanceFusion module, as shown in Algorithm 1, to achieve cross-modal alignment of instance features. For positive samples, we project LiDAR instance features onto the camera BEV space and consider the matched camera instance features as positive samples. For negative samples, we select the $K$ nearest neighboring instance features around the camera positive samples as negative samples. During inference, LiDAR instance features calculate similarity scores with $K$ neighboring camera instances and select the instance feature with the highest score as the matching aligned instance feature. The whole process of InstanceFusion module are as follows. To construct positive and negative sample pairs for contrastive learning, we establish neighboring instance feature relationships based on LiDAR and camera Instance features using a graph matching approach. During training, LiDAR instance features are projected onto the camera BEV space. Instances, that overlap each other, are directly considered positive samples, while those with close proximity in the KNN graphs are considered negative samples. During the inference process, the LiDAR instance features calculate the similarity score between the $K$ camera instances through graph matching, and the camera instance feature with the highest score is selected as the matching instance.
@@
First, due to the absence of depth errors in the LiDAR BEV feature implementation, we utilize the generated LiDAR RoI features to retrieve the corresponding camera RoI features. For any pair of bounding box predictions from $P_{C}$ with $N$ samples and $P_{L}$ with $M$ samples, we calculate the Intersection over Union (IoU) by computing the intersection and union areas of their 2D bounding boxes. We define the bounding box sets $\text{B}_{C}^{RoI}=\{(x_{i},y_{i},w_{i},h_{i})|i\in[1,2...N]\}$ and $\text{B}_{L}^{RoI}=\{(x_{i},y_{i},w_{i},h_{i})|i\in[1,2...M]\}$ based on the positions $(\delta_{x},\delta_{y})$ and dimensions $(w,h)$ of proposals. To obtain the groups of positive samples for $F^{pos}$, we execute that
@@
After obtaining the positive samples, we proceed to construct negative samples. Due to the impact of feature misalignment, offsets occur in a region. We utilize graph matching like the KNN algorithm to construct the neighbors’ relations of positive samples and consider each positive sample’s neighbor features as negative. Specifically, the construction of negative samples is as follows.
@@
V Experiments
In this section, we present the experimental setup of ContrastAlign and evaluate the performance of 3D object detection on the nuScenes [51] dataset. Additionally, we follow the solution provided by Ref. [52] to simulate feature misalignment scenarios.
TABLE I: Comparison with the SOTA methods on the nuScenes validation set. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘ T.C.’ are short for construction vehicles, motorcycles, pedestrians, and traffic cones.
TransFusion-L [28] L 65.1 70.1 86.5 59.6 25.4 74.4 42.2 74.1 72.1 56.0 86.6 74.1